# Orthogonal Block Kaczmarz Algorithm Based on Preprocessing Technology

### v3

@@ 3.2 The Proposed methods
As mentioned above, both RBK(k) and GREK(k) algorithms use the k-means algorithm to group matrix rows. Undoubtedly, k-means clustering is one of the most effective methods for dividing data into different groups. However, due to only considering the computational efficiency after partitioning and ignoring the computational cost of partitioning, the overall computational efficiency of the classic block Kaczmarz algorithm is not high. Therefore, to address these problems, a simple and fast block Kaczmarz algorithm that combines preprocessing techniques and orthogonal partitioning ideas is proposed in this paper.
Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$ as follows.
$PAP^{T}$ (3.1)
where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements of the original coefficient matrix $A$ closer to the main diagonal, i.e., it makes the structure of matrix $A$ closer to the banded matrix form, and $P$ is a substitution matrix (i.e., there is $P^{T}P=E$), so the linear system (1.1) becomes
@@
Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$. Since $\widetilde{A}$ is an approximate banded matrix, then it is sufficient to make the row blocks orthogonal to each other by doing appropriate row chunking of $\widetilde{A}$, see Figure 2.
Due to the natural orthogonality of banded matrices (see Figure 2), we directly chunk $\widetilde{A}$ uniformly in order. Suppose that the matrix $\widetilde{A}$ is divided into $k$ blocks, the first $k-1$ blocks take $\left[m/k\right]$ elements each in order, and the remaining elements are stored in the $k$-th block, and the corresponding chunking is done for $f$,
@@
As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution. Then, according to the analysis in Section 3.1,, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non-orthogonal hyperplanes, see Figure 1. So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
@@
Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{i}$ as $\overline{A}_{i}$, $i=1,2,...,k$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{i},\overline{A}_{j}\right)$, the two-paradigm number of $\overline{A}_{i}$ and $\overline{A}_{j}$, and then compute the cosine of the angle between the centroid coordinates according to the angle cosine formula $C\left(i,j\right)=\frac{G\left(i,j\right)}{\left\|\overline{A}_{i}\right\|_{2}\left\|\overline{A}_{j}\right\|_{2}}$, which constitutes a table $C$ of the cosine values, and $C\left(i,j\right)$ is used to measure orthogonality between blocks ${A}_{i}$ and blocks ${A}_{j}$. And since $\widetilde{A}$ is an approximate banded matrix, it is reasonable and convenient to use the cosine of the angle between the centroid coordinates to reflect the orthogonality between row blocks. And we divide the matrix $\widetilde{A}$ into $k$ blocks beforehand, and we only need to calculate the cosine of the angle between these $k$ blocks, and since $C$ is a symmetric matrix, we only need to calculate $\frac{k\left(k-1\right)}{2}$ times in practice. Since $k\ll m$, this is a fairly small number.
@@
Next, based on our analysis above, we aim to accelerate the Kaczmarz method by separating the two row blocks that are nearly orthogonal, i.e., $C(i,j)=0$, into one class. Here we set a threshold $thr$ such that when $C(i,j)<thr$, block $A_{\tau_{i}}$ and block $A_{\tau_{j}}$ are considered orthogonal, i.e., so that $C(i,j)=0$. However, it is obvious that for any block $A_{\tau_{i}}$, there is most likely more than one block that is orthogonal to it, and this can easily result in the situation where some rows blocks are in several classes at the same time. In order to avoid this situation, we stipulate that for each block $A_{i}$, $i=1,2,..k$, the first block $A_{j}\in\left\{A_{i},A_{i+1},...,A_{k}\right\}$ orthogonal to $A_{i}$ is selected and placed in a class, noting the current class as $class\left\{i\right\}$, and requiring that there is no duplication of blocks between classes. We merge these orthogonal classes, denoted $Oclass$. For the blocks that do not go into $Oclass$, in order to avoid the non-convergence situation caused by the loss of information, we assign all of these blocks in a new class, denoted $Nclass$. This regulation is reasonable because of the nature of the A-approximate banded matrix. The solution is obtained by the block Kaczmarz iterative method on both categories:
$\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right)$ (3.3)
@@
Finally, the solution of the original system is obtained as $x=P^{T}\widetilde{x}$.
Algorithm 5 POBK Algorithm
@@
3:Preprocess the original linear system by RCM, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
@@
5:Compute the cosine of the angle $C\left(i,j\right)$ between the centroid coordinates, and constitutes a table $C$ of the cosine values.
6:According to the above regulations, place all two-by-two mutually orthogonal blocks in a category $Oclass$, place all remaining blocks in a new class $Nclass$.
@@
Unlike the GRBK, RBK(k) and GREBK(k) methods, our proposed method does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals. In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal. Our proposed idea of performing $\frac{k\left(k-1\right)}{2}$ calculations of the angle cosine values and then orthogonal chunking avoids this problem to a large extent.On the one hand, the computation time of the angle cosine value is very short because $k\ll m$. On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
4 Convergence analysis
Assuming that the exact solution of the original linear system $Ax=f$ is $x_{*}$, and the exact solution of the preprocessed linear system $\widetilde{A}\widetilde{x}=\widetilde{f}$ is $\widetilde{x}_{*}$, where $P$ is the permutation matrix, then
@@ 4 Convergence analysis
It can be seen that the convergence analysis of the original linear system and the preprocessed linear system is consistent, therefore, we will not differentiate this in the future.
Lemma 1 [25]. Assumes that the singular value decomposition is performed on the submatrix $A_{i}\in{{\mathbb{R}}^{m_{i}\times m}}$,
$A_{i}=U_{i}\Sigma_{i}V_{i}^{H}$ (4.2)
where $U_{i}$ is a $m_{i}\times m_{i}$ unitary matrix. $\Sigma_{i}$ is a $m_{i}\times m$ diagonal matrix. $V_{i}^{H}$, the conjugate transpose of $V_{i}$, is a unitary matrix of $m\times m$ order. The elements on the diagonal of matrix $\Sigma_{i}$ are singular values of $A_{i}$. Then the Moore–Penrose pseudoinverse of $A_{i}$ $A_{i}^{\dagger}=V_{i}\Sigma_{i}^{H}U_{i}^{H}$, and $A_{i}^{\dagger}A_{i}=V_{i}MV_{i}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{i}$.
Lemma 2 [13]. For any $A\in{{\mathbb{R}}^{n\times m}}$, we have
$\sigma_{min}(A^{\dagger})=\frac{1}{\sigma_{max}(A)},$
where $\sigma_{min}\left(.\right)$ denotes the smallest nonzero singular value of a matrix and $\sigma_{max}\left(.\right)$ denotes the largest singular value of a matrix.
Similar to Ref. [16], we have the following convergence theorem.
Theorem 1. If the linear system (1.1) $Ax=f$ is consistent, then the POBK algorithm produces a sequence of iterations $\left\{x_{j}\right\}_{j=0}^{\infty}$ that converges in expectation to a unique minimum norm solution $x_{\star}=A^{\dagger}f$. In particular, we have the following linear convergence rate in expectation:
@@
Proof. When $\tau_{j}\in Nclass$, the POBK algorithm involves one block per iteration and computes the linear system by the following form
@@
According to Lemma 1, there are $A_{\tau_{j}}=U_{\tau_{j}}\Sigma_{\tau_{j}}V_{\tau_{j}}^{H}$, $A_{\tau_{j}}^{\dagger}=V_{\tau_{j}}\Sigma_{\tau_{j}}^{H}U_{\tau_{j}}^{H}$, and therefore $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}=V_{\tau_{j}}MV_{\tau_{j}}^{H}$. Thus $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}$ is an orthogonal projection operator, then
$x_{j+1}-x_{*}$ $=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)-x_{*}$ (4.5)
$=x_{j}-x_{*}+A_{\tau_{j}}^{\dagger}\left(A_{\tau_{j}}x_{*}-A_{\tau_{j}}x_{j}\right)$
$=\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right).$
@@
$\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$ (4.6)
$=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$
$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}.$
@@
$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]$ $=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\mathbb{E}_{j}\left[\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$ (4.7)
$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sum_{\tau\in\left\{1,2,...,k\right\}}p_{\tau}\left[\left\|A_{\tau}^{\dagger}A_{\tau}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$
$\leqslant\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sigma_{min}^{2}\left(A_{\tau}^{\dagger}\right)\sigma_{min}^{2}\left(A_{\tau}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}$
$\leqslant\left(1-\frac{\sigma_{min}^{2}(A_{\tau})}{\sigma_{max}^{2}(A_{\tau})}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2},$
and assuming that $\lambda_{min}^{\tau}$=$\sigma_{min}^{2}\left(A_{\tau}\right)$, $\lambda_{max}^{\tau}$=$\sigma_{max}^{2}\left(A_{\tau}\right)$. Then we have
@@
When $\tau_{1},\tau_{2}\in Oclass$, each iteration of the POBK algorithm involves two blocks, that is, iterates once on the basis of the original iteration, then obviously,
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{(j+1)/2}-x_{*}\right\|_{2}^{2}$ (4.10)
$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{2}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
@@
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{1}}\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{2}}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}$ (4.11)
$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
@@
Remark 1. As can be seen from the above proof, the POBK algorithm converges as expected and the convergence rate is related to the maximum singular value of the block, the minimum non-zero singular value, and the number of chunks. The power of the convergence rate of the POBK algorithm is much larger than the power of the general Kaczmarz algorithm, and theoretically there will be a much faster convergence speed. In order to visualize the convergence speed of the POBK algorithm more intuitively, below we explain the reason why the POBK algorithm converges faster from the perspective of the hyperplane pinch angle.
@@
$cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}}$ (4.12)
@@
From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the value of the cosine of the hyperplane pinch angle. The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm. On the contrary, the smaller the angle $\theta$ between the hyperplanes (closer to $0^{\circ}$), the closer the value of $cos\theta$ is to 1, and the slower the convergence speed of the algorithm. Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between two blocks that are mutually orthogonal in $Oclass$ is close to $90^{\circ}$, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms. Therefore, the iterative solution will approach the exact solution more quickly. Although the blocks in $Nclass$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate banded matrix, the orthogonality between the blocks in the matrix $\widetilde{A}$ is better than the original matrix $A$. Meanwhile, two nearly orthogonal blocks in $Oclass$ have a high probability of being non-adjacent, then the matrices in $Nclass$, although not orthogonal to each other, will have a much lower probability of being adjacent to each other, i.e., the number of times to compute the cosine value of the smaller pinch angle will be much lower, and according to Equation (4.14), its convergence speed is also faster than that of the GRBK, RBK (k), GREBK (k) and aRBK algorithms.
5 Numerical experiments
5.1 Numerical simulation
In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK(k), GREBK(k), aRBK and POBK methods, verifying the effectiveness and efficiency of the algorithm. The matrices used for numerical experimental tests are 20 sparse matrices selected from the $SuiteSparse\;Matrix\;$ $Collection$ with application background and varying in size, condition number and density. Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the linear system. We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector space.

### v4

@@ 3.3 The Proposed methods
As mentioned above, both RBK(k) and GREK(k) algorithms use the K-means algorithm to group matrix rows. Undoubtedly, K-means clustering is one of the most effective methods for dividing data into different groups. However, due to only considering the computational efficiency after partitioning and ignoring the computational cost of partitioning, the overall computational efficiency of the classic block Kaczmarz algorithm is not high. Therefore, in order to solve these problems, this paper proposes a simple and fast block Kaczmarz algorithm that combines matrix reordering techniques and orthogonal partitioning ideas.
Based on the above analysis, We reorder the original matrix $A$ as follows:
$PAP^{T},$ (3.1)
where $P$ is obtained by the RCM reordering algorithm and $P$ is a substitution matrix (i.e., $P^{T}P=E$), so the linear system (1.1) becomes
@@
Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$. Since $\widetilde{A}$ is an approximate banded matrix, then it is sufficient to make the row blocks orthogonal to each other by doing appropriate row blocking of $\widetilde{A}$, see Figure 3.
Figure 3:   The orthogonality representation of the blckhole matrix after reordering with the RCM algorithm.
Due to the natural orthogonality of banded matrices, we directly chunk $\widetilde{A}$ uniformly in order. Suppose that the matrix $\widetilde{A}$ is divided into $k$ blocks, the first $k-1$ blocks take $\left[m/k\right]$ elements each in order, and the remaining elements are stored in the $k$-th block, and the corresponding blocking is done for $\widetilde{f}$,
@@
As is well known, the Kaczmarz method iteratively projects the current estimate orthogonally onto successive hyperplanes to approach the exact solution. As analyzed in Section 3.1, orthogonal hyperplanes enable accelerated convergence: pairwise orthogonal projections (e.g., $H_{1}\bot H_{3}$ in Figure 1) achieve quadratic convergence within two iterations, while non-orthogonal configurations require significantly more steps. This motivates a critical innovation: grouping mutually orthogonal hyperplanes into block projections to eliminate redundant iterations and accelerate convergence—a strategy central to our proposed method.
@@
Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{i}$ as $\overline{A}_{i}$, $i=1,2,...,k$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{i},\overline{A}_{j}\right)$, the two-paradigm number of $\overline{A}_{i}$ and $\overline{A}_{j}$, and then compute the cosine of the angle between the centroid coordinates according to the angle cosine formula $cos\theta_{ij}=\frac{G\left(i,j\right)}{\left\|\overline{A}_{i}\right\|_{2}\left\|\overline{A}_{j}\right\|_{2}}$, which constitutes a table $C$ of the cosine values, and $cos\theta_{ij}$ is used to measure orthogonality between blocks ${A}_{i}$ and blocks ${A}_{j}$. And since $\widetilde{A}$ is an approximate banded matrix, it is reasonable and convenient to use the cosine of the angle between the centroid coordinates to reflect the orthogonality between row blocks. And we divide the matrix $\widetilde{A}$ into $k$ blocks beforehand, and we only need to calculate the cosine of the angle between these $k$ blocks, and since $C$ is a symmetric matrix, we only need to calculate $\frac{k\left(k-1\right)}{2}$ times in practice. Since $k\ll m$, this is a fairly small number.
@@
Next, based on our analysis above, we aim to accelerate the Kaczmarz method by separating the two row blocks that are nearly orthogonal, i.e., $cos\theta_{ij}\approx 0$, into one class. Here we set a threshold $thr$ such that when $cos\theta_{ij}<thr$, block $A_{\tau_{i}}$ and block $A_{\tau_{j}}$ are considered orthogonal. However, it is obvious that for any block $A_{\tau_{i}}$, there is most likely more than one block that is orthogonal to it, and this can easily result in the situation where some rows blocks are in several classes at the same time. In order to avoid this situation, we stipulate that for each block $A_{i}$, $i=1,2,..k$, the first block $A_{j}\in\left\{A_{i},A_{i+1},...,A_{k}\right\}$ orthogonal to $A_{i}$ is selected and placed in a class, noting the current class as $class\left\{i\right\}$, and requiring that there is no duplication of blocks between classes. We merge these orthogonal classes, denoted $Oclass$. For the blocks that do not go into $Oclass$, in order to avoid the non-convergence situation caused by the loss of information, we assign all of these blocks in a new class, denoted $Nclass$. This regulation is reasonable because of the nature of the A-approximate banded matrix. The solution is obtained by the block Kaczmarz method on both categories:
$\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right),$ (3.3)
@@
Finally, the solution of the original system is obtained as $x=P^{T}\widetilde{x}$. See Algorithm 4 for details.
Algorithm 4 OBK-RCM Algorithm
1:$A$, $f$, $x_{0}$, $k$, $thr$, $l$.
@@
3:Reordering matrices using the RCM algorithm, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
@@
5:Compute the cosine of the angle $cos\theta_{ij}$ between the centroid coordinates, and constitutes a table $C$ of the cosine values.
6:Chunking $Oclass$ and $Nclass$ according to the algorithm description.
@@
Unlike the RBK, RBK(k) and GREBK(k) methods, the OBK-RCM algorithm does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals. In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal. Our proposed idea of performing $\frac{k\left(k-1\right)}{2}$ calculations of the angle cosine values and then orthogonal chunking avoids this problem to a large extent.On the one hand, the computation time of the angle cosine value is very short because $k\ll m$. On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
3.4 Convergence analysis
Assuming that the exact solution of the original linear system $Ax=f$ is $x_{*}$, and the exact solution of the reordered linear system $\widetilde{A}\widetilde{x}=\widetilde{f}$ is $\widetilde{x}_{*}$, where $P$ is the permutation matrix, then
@@ 3.4 Convergence analysis
It can be seen that the convergence analysis of the original linear system and the reordered linear system is consistent, therefore, we will not differentiate this in the future.
Theorem 1. If the linear system (1.1) $Ax=f$ is consistent, then the OBK-RCM algorithm produces a sequence of iterations $\left\{x_{j}\right\}_{j=0}^{\infty}$ that converges in expectation to a unique minimum norm solution $x_{\star}=A^{\dagger}f$. In particular, we have the following linear convergence rate in expectation:
@@
Proof. When $\tau_{j}\in Nclass$, the OBK-RCM algorithm involves one block per iteration and computes the linear system by the following form
@@
According to the singular value decomposition, there are $A_{\tau_{j}}=U_{\tau_{j}}\Sigma_{\tau_{j}}V_{\tau_{j}}^{H}$, $A_{\tau_{j}}^{\dagger}=V_{\tau_{j}}\Sigma_{\tau_{j}}^{H}U_{\tau_{j}}^{H}$, and therefore $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}=V_{\tau_{j}}MV_{\tau_{j}}^{H}$. Thus $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}$ is an orthogonal projection operator, then
$\begin{split}x_{j+1}-x_{*}&=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)-x_{*}\\ &=x_{j}-x_{*}+A_{\tau_{j}}^{\dagger}\left(A_{\tau_{j}}x_{*}-A_{\tau_{j}}x_{j}\right)\\ &=\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right).\end{split}$ (3.7)
@@
$\begin{split}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}&=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)\\ &=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)\\ &=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}.\end{split}$ (3.8)
@@
$\begin{split}\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]&=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\mathbb{E}_{j}\left[\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]\\ &=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sum_{\tau\in\left\{1,2,...,k\right\}}p_{\tau}\left[\left\|A_{\tau}^{\dagger}A_{\tau}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]\\ &\leqslant\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sigma_{min}^{2}\left(A_{\tau}^{\dagger}\right)\sigma_{min}^{2}\left(A_{\tau}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}\\ &\leqslant\left(1-\frac{\sigma_{min}^{2}(A_{\tau})}{\sigma_{max}^{2}(A_{\tau})}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2},\end{split}$ (3.9)
where $\sigma_{min}(.)$ denotes the minimum non-zero singular value of a matrix, $\sigma_{max}(.)$ denotes the maximum singular value of a matrix. It may be useful to set $\lambda_{min}^{\tau}$=$\sigma_{min}^{2}\left(A_{\tau}\right)$, $\lambda_{max}^{\tau}$=$\sigma_{max}^{2}\left(A_{\tau}\right)$. Then we have
@@
When $\tau_{1},\tau_{2}\in Oclass$, each iteration of the OBK-RCM algorithm involves two blocks, that is, iterates once on the basis of the original iteration, then obviously,
$\begin{split}\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}&\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{(j+1)/2}-x_{*}\right\|_{2}^{2}\\ &\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{2}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.\end{split}$ (3.12)
@@
$\begin{split}\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}&\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{1}}\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{2}}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}\\ &\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.\end{split}$ (3.13)
@@
Remark 1. From the above proof, it can be seen that the OBK-RCM algorithm converges as expected, and the rate of convergence is closely related to the maximum singular value of the block, the minimum non-zero singular value, and the number of chunks. When the submatrix orthogonality is enhanced (i.e., $\lambda_{max}^{\tau}\approx\lambda_{min}^{\tau}$), the ratio tends to 1 and the rate of convergence increases significantly. This explains the key role of orthogonal chunking of the matrix after reordering in accelerating the convergence.
According to the analysis in Section 3.1, the influence of the angle between the matrix row blocks on the convergence speed is crucial. In order to visualize the convergence speed of the OBK-RCM algorithm more intuitively, the following section explains the reason why the OBK-RCM algorithm converges faster from the perspective of the angle between hyperplanes.
@@
$cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}},$ (3.14)
@@
From Equation (3.16), it can be seen that the convergence speed of the Kaczmarz algorithm is closely related to the cosine of the hyperplane angle. The larger the hyperplane angle $\theta$ is (the closer to $90^{\circ}$), and the closer $cos\theta$ is to 0, the faster the algorithm converges; on the other hand, the smaller the hyperplane angle $\theta$ is (the closer to $0^{\circ}$), and the closer $cos\theta$ is to 1, the slower the convergence speed of the algorithm is. The OBK-RCM algorithm utilizes the RCM sorting algorithm to reorder the matrices so that the hyperplanes $H_{1}$ and $H_{2}$ are approximately orthogonal ($cos\theta\approx 0$), and the error is rapidly decayed so that the convergence speed is rapidly improved.
Briefly, the OBK-RCM algorithm achieves superior convergence over traditional block Kaczmarz methods (RBK, RBK(k), GREBK(k), aRBK) through two synergistic mechanisms:
  1. 1.
    RCM-Driven Orthogonality Enhancement: By reordering the matrix into a banded structure (Table 1), OBK-RCM forces inter-block angles toward orthogonality, enabling exact solution convergence in two iterations for orthogonal hyperplanes (Figure 1), versus polynomial-time scaling for non-orthogonal cases.
  2. 2.
    Intelligent Block Scheduling: The $Oclass$/$Nclass$ partitioning (Section 3.3) systematically minimizes adjacency of non-orthogonal blocks in $Nclass$. Although $Nclass$ blocks lack strict orthogonality, their residual angular cosines remain significantly lower than in the original matrix $A$ (Equation 3.16), as RCM reordering inherently suppresses small-angle configurations (Figure 2).
This dual strategy–orthogonal acceleration in $Oclass$ and mitigated error propagation in $Nclass$–collectively ensures OBK-RCM’s 10-50 times speedup over baseline methods, as quantified in Table 4.
3.5 Numerical experiments
In this section, we conducted some experiments to compare the convergence speed of RBK, RBK(k), GREBK(k), aRBK and OBK-RCM methods, verifying the effectiveness and efficiency of the algorithm. The matrices used for numerical experimental tests are 15 sparse matrices selected from the $SuiteSparse\;Matrix\;$ $Collection$ with application background and varying in size, condition number and density. Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the linear system. We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector space.